Steve is half as old as Stacey. The sum of their ages five years ago was 8. How old are Steve and Stacey now?
Answer by nyc_function(2741)

Steve's age now = x/2
Stacey's age now = x
Steve's age 5 years ago = (x/2) - 5
Stacey's age 5 years ago = x - 5
Here is your equation:
x + (x/2) - 10 = 8
Can you finish now?